Marc Horowitz 21.512

Uses of Body Symbolism

The human body is a form on which many symbols are based. We use our own bodies to symbolize many things, and may describe other symbols in terms of the body.

One common form of symbolism relating to the body is adornment. People decorate their bodies as an external representation of some aspect of the individual. Appearance can symbolize occupation, age, wealth, social position, marriage status, or many other attributes within a society.

Within our society, there are several kinds of adornment which are used symbolically. The most obvious example of this is clothing. Clothing can be symbolic of the occupation of the person wearing it. Many occupations, such as policemen, doctors, judges, and clergymen, all wear distinctive clothing. Our society makes sure that people in important and honorable occupations are distinguished by their dress. Clothing can also denote wealth. Someone who wants to be recognized as wealthy will wear expensive clothing. Certain materials and styles are symbolic of wealth. Styles of clothing can also represent ethnicity or religion. People from different regions will often have different styles of clothing specific to their region. Clothing symbols can often come into play in other circumstances than someone wearing the clothes. A cartoon with a picture of a pig wearing a police badge and hat would be symbolic of a negative view of policeman. This uses the body symbolism of the uniform together with the animal symbolism of a pig as a dirty animal together to make a societal comment about the people who normally wear the policeman’s uniform.

Another example of body adornment symbolism is hairstyle. Across many societies, hairstyle is a common symbol of social order. In Western society, hairstyle often depicts the personality of the individual. In general, men wear their hair short, and women wear it long. But a man or woman who wants to appear different may do the opposite of the norm. A long-haired man is a “hippie” or some other non-mainstream subclass. A short-haired woman is often considered “masculine”. Hair color can also symbolize an attitude towards life. “Blondes has more fun” is a good example of this. Shaving the head has a significant symbolic value. It represents a break from both the short-haired conservatives and the long-haired liberals.

In other societies, hair symbolizes other things. In the Tchikrin tribe, hair is a symbol of sexual maturity. Both women and men grow their hair long once they have proven themselves sexually by having children. Until then, their hair is kept short. The Tikopia are exactly opposite our own traditional system: Men wear long hair, and women wear short hair. Where we consider short hair to be symbolic of mainstream culture and power, they have the same view of long hair.

Another form of body symbolism is metaphorical. It is common to hear various parts of the body used to describe another system.

We use body human as a symbol for the body politic. We might speak of civil servants as the hands of the government, as they are the people who get day-to-day work done. Newspaper reporters are the “eyes and ears” of society, watching everything which goes on. A strong economic heartbeat is crucial to survival. Crime is a disease, one which needs to be found and cured before it can cause great harm to the body.

Other examples are not hard to find. The computer controlling a machine is often referred to as the brains of the system. A fast, powerful car has a lot of muscle. A lever arm is a useful tool. A chair stands on its legs. There are a whole body of such examples.

Body symbolism is common to almost all societies. Although it is viewed differently by different groups of people, it is a common form which can be interpreted in many interesting ways. This makes it a rich source of symbolism of all kinds.